# moonboot: An R Package Implementing m-out-of-n Bootstrap Methods

### v2

@@ 4   Results
For all four estimators, we have evaluated the methods for estimating $\tau_{n}$ and the different choices for $m$ in the “n-choose-m” bootstrap by means of Monte Carlo simulations. We have computed confidence intervals with a nominal coverage probability $P_{cov}$ of $0.95$ and estimated the actual coverage probability by repeating the computation $N=10^{4}$ times. This means that the accuracy (i.e. the width of a 95% confidence interval) of the estimated $P_{cov}$ is about ±$\pm 0.005$. Using a larger $N$ would have made the simulation for the method by Goetze intractable. All simulations were done with $R=1000$ bootstrap repetitions.
@@ 4.1   Estimation of $\tau_{n}$
method estimator true $\tau_{n}$ $(\beta_{1},\beta_{5})$ $n$ estimated $\tau_{n}$ $n$ estimated $\tau_{n}$
@@
Table 2: Dependency of the estimation of $\tau_{n}$ on the chosen endpoints for the sequence $m_{i}=n^{\beta_{i}}$ and two different values for $n$. The better choice is marked in bold for each estimator. The estimated values for $\tau_{n}$ have been averaged over 100 estimations.
Our simulations confirmed that the choice for the extreme values of the trial values $m_{i}=n^{\beta_{i}}$ for $i=1,\ldots,5$ indeed influence the accuracy of the estimation of $\tau_{n}$. For the mean estimator, e.g., we achieved the best results for $\beta_{1}=0.2$ and $\beta_{5}=0.5$, whereas this choice was poor for xicor and max (see Tbl. 2). The results for both estimations, by means of the variance or quantile ranges, are very similar. We thus conclude that it does not matter which method is used. The results for xicor, however, are a warning that the estimation of $\tau_{n}$ with the method by Bertail et al. (1999) can be grossly inaccurate, unless the range of the $\beta_{i}$ has been luckily guessed.
@@
Figure 3: Coverage probability $P_{cov}$ of the n-choose-m bootstrap for the choices $m=n^{\beta}$ with $\beta\in\{\frac{1}{2},\frac{1}{3},\frac{3}{4}\}$. Note that the $n$-axis uses a logarithmic scale.
@@ 4.2   Choice of $m$
Figure 4: Comparison of the three data based methods for choosing $m$ in the n-choose-m boostrap. The upper plots show the mean value $m(n)$ together with error bars ±$\pm\operatorname{sd}(m)$ and a least squares fit $\log(m)=\log(cn^{\beta})$. The bottom plots show the resulting coverage probabilities. Note that the $n$-axis uses a logarithmic scale.
The results for the data based methods for choosing $m$ are summarized in Fig. 4. The method by Politis et al. (1999) cannot be recommended, because the resulting coverage probability never approached the nominal value in our simulations. For the shorth, it was only about 0.7 even for high values of $n$. For the other estimators, it was somewhat greater, but nevertheless too small. This behavior can be understood from the curves showing $m(n)$ in Fig. 4: The method by Politis et al. (1999) always chooses an $m$ proportional to $n$, which means that the necessary condition $\lim_{n\to\infty}m/n=0$ is violated.
@@
Figure 5: Comparison of the n-choose-m bootstrap with the ordinary n-out-of-n (basic) bootstrap for the bootstrap consistent estimator mean. Note that the $n$-axis uses a logarithmic scale.
Out of curiosity, we have also compared the n-choose-m bootstrap with with the ordinary n-out-of-n bootstrap for the estimator mean, which is bootstrap consistent and thus allows for a comparison. As can be seen in Fig. 5, the m-out-of-n bootstrap is competitive with the basic bootstrap for large $n\gtrsim 2000$ and a luckily chosen $m$. For a data based choice of $m$ with the method by Bickel and Sakov (2008), the coverage probability is smaller, although the difference in the interval length is small. This demonstrates that, for bootstrap consistent estimators, the usual n-out-of-n bootstrap is preferable, because the best choice for $m$ is not known, in general.

### v3

@@ 4  Results
For all four estimators, we have evaluated the methods for estimating $\tau_{n}$ and the different choices for $m$ in the m-out-of-n bootstrap without replacement by means of Monte Carlo simulations. We have computed confidence intervals with a nominal coverage probability $P_{cov}$ of $0.95$ and estimated the actual coverage probability by repeating the computation $N=10^{4}$ times. This means that the accuracy (i.e. the width of a 95% confidence interval) of the estimated $P_{cov}$ is about $\pm 0.005$. Using a larger $N$ would have made the simulation for the method by Goetze intractable. All simulations were done with $R=1000$ bootstrap repetitions.
@@ 4.1  Estimation of $\tau_{n}$
method estimator true $\beta$ $(\gamma_{1},\gamma_{5})$ $n$ estimated $\beta$ $n$ estimated $\beta$
@@
Table 2: Dependency of the estimation of $\tau_{n}=n^{\beta}$ on the chosen endpoints for the sequence $m_{i}=n^{\gamma_{i}}$ and two different values for $n$. The better choice is marked in bold for each estimator. The estimated values for the power $\beta$ in $\tau_{n}$ have been averaged over 100 estimations.
@@
Figure 3: Coverage probability $P_{cov}$ of the m-out-of-n bootstrap without replacement for the choices $m=n^{\beta}$ with $\beta\in\{\frac{1}{2},\frac{1}{3},\frac{3}{4}\}$. Note that the $n$-axis uses a logarithmic scale.
Our simulations confirmed that the choice for the extreme values of the trial values $m_{i}=n^{\gamma_{i}}$ for $i=1,\ldots,5$ indeed influence the accuracy of the estimation of $\tau_{n}$. For the mean estimator, e.g., we achieved the best results for $\gamma_{1}=0.2$ and $\gamma_{5}=0.5$, whereas this choice was poor for xicor and max (see Tbl. 2). The results for estimation by means of the variance or quantile ranges are similar. We thus conclude that it does not matter which method is used. The results for xicor, however, are a warning that the estimation of $\tau_{n}$ with the method by Bertail et al. (1999) can be grossly inaccurate, unless the range of the $\gamma_{i}$ has been luckily guessed.
@@ 4.2  Choice of $m$
Figure 4: Comparison of the three data based methods for choosing $m$ in the m-out-of-n bootstrap without replacement. The upper plots show the mean value $m(n)$ together with error bars $\pm\operatorname{sd}(m)$ and a least squares fit $\log(m)=\log(cn^{\beta})$. The bottom plots show the resulting coverage probabilities. Note that the $n$-axis uses a logarithmic scale.
The results for the data based methods for choosing $m$ are summarized in Fig. 4. The method by Politis et al. (1999) cannot be recommended, because the resulting coverage probability never approached the nominal value in our simulations. For the shorth, it was only about 0.7 even for high values of $n$. For the other estimators, it was somewhat greater, but nevertheless too small. This behavior can be understood from the curves showing $m(n)$ in Fig. 4: The method by Politis et al. (1999) always chooses an $m$ proportional to $n$, which means that the condition $\lim_{n\to\infty}m/n=0$ is violated.
@@
Figure 5: Comparison of the m-out-of-n bootstrap without replacement with the ordinary n-out-of-n (basic) bootstrap for the bootstrap consistent estimator mean. Note that the $n$-axis uses a logarithmic scale.
Out of curiosity, we have also compared the m-out-of-n bootstrap without replacement with the ordinary n-out-of-n bootstrap for the estimator mean, which is bootstrap consistent and thus allows for a comparison. For bootstrap consistent estimators with additional smoothness properties [^6] Most notably that it is twice Fréchet differentiable., it was generally proven by Bickel et al. (1997) that estimates based on the m-out-of-n bootstrap with or without replacement are less efficient than those based on the n-out-of-n bootstrap. The mean is an example for such an estimator. As can be seen in Fig. 5, the m-out-of-n bootstrap is competitive with the basic bootstrap only for large $n\gtrsim 2000$ and a luckily chosen $m$. For a data based choice of $m$ with the method by Bickel and Sakov (2008), the coverage probability is smaller, although the difference in the interval length is small. This again demonstrates that, for bootstrap consistent estimators, the usual n-out-of-n bootstrap is preferable, especially because the best choice for $m$ is not known, in general.
5  Violation of the assumptions
The assumptions (1) & (2), under which the m-out-of-n bootstrap works, are quite weak, but there are nevertheless some estimators that violate these conditions. For example, inconsistent estimators, i.e., estimators that do not converge in probability to the true parameter value, violate the assumptions because consistency is a necessary condition for the assumptions to hold, as shown in section 2. Using moonboot in such a case is an error on the side of the user, and if the user has determined the scaling factor $\tau_{n}$ by an analysis of the estimator, he usually will be aware of this. It might be, however, that a user relies on the fully automated estimation of $\tau_{n}$ provided by moonboot and thus is not aware of a violation of the assumptions.
We therefore have tested our package with two unbiased, yet inconsistent estimators, too. Please note that these estimators are bizarre examples and are not meant to be used in practice. The first estimator is only the very first observation as an estimator for the mean $\mu$ of the unsymmetric distribution moonboot::dpower(..,2,0,1):
$\hat{\mu}_{1}=X_{1}$ (10)
The distribution of this estimator is the same for all sample sizes, and the scaling factor is thus $\tau_{n}\propto 1$, which fulfills condition (2), but violates condition (1). The second estimator estimates the parameter $\lambda$ of a Poisson distribution. For a Poisson distribution, both mean and variance are $\lambda$, which allows for the construction of an unbiased estimator for $\lambda$ from the unbiased estimators $\overline{X}$ for the mean and $S^{2}$ for the variance as follows:
$\hat{\lambda}_{n}=n\overline{X}-(n-1)S^{2}=\sum_{i=1}^{n}X_{i}-\sum_{i=1}^{n}\left(X_{i}-\overline{X}\right)^{2}$ (11)
As both $\overline{X}$ and $S^{2}$ are root-$n$ consistent estimators, the scaling factor for condition (2) to hold is $\tau_{n}\propto n^{-1/2}$, which again is in violation of condition (1).
As the assumptions (1) & (2) are only sufficient, but not necessary conditions for the m-out-of-n bootstrap, we first estimated the coverage probability of the “basic” m-out-of-n bootstrap interval when the correct scaling factor $\tau_{n}$ is provided. For $\hat{\mu}_{1}$, the coverage probability obviously neither depends on $n$ nor on the subsampling size $m$, and it turned out to be about $0.52$ for a nominal $0.95$ interval. For $\hat{\lambda}_{n}$, the coverage probability decreased with increasing $n$ and was only about $0.17$ for $n=5000$ and $m=\sqrt{n}$. These results show that the m-out-of-n bootstrap indeed does not work for these inconsistent estimators.
We then estimated the scaling factors with estimate.tau() for $n=100$ and $n=500$, and, for all parameter settings, the mean estimates for $\beta$ in $\tau_{n}=n^{\beta}$ were less than $0.002$ for $\hat{\mu}_{1}$ and less than 0.40-0.40 for $\hat{\lambda}_{n}$. This means that, for these estimators, the violation of the assumptions can be automatically detected due to suspiciously small estimated values for $\beta$. In order to make the user aware of this issue, mboot.ci() gives a warning if $\tau_{n}$ decreases or if its increase is suspiciously small, i.e., if it increases at a slower rate than $n^{0.01}$.
@@ 6  Conclusions
Acknowledgements
We thank the anonymous reviewers for their valuable comments and for the suggestion to investigate the case when the assumptions are violated.